OpenStudy (anonymous):

Integrate. How do I integrate (e^2x)/(1+e^x)?

OpenStudy (callisto):

Is this the question? \[\int\frac{e^{2x}}{1+e^x}dx\]

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Let u = 1+e^x, that is e^x=u-1 du = e^x dx, so we get \[\int\frac{e^{2x}}{1+e^x}dx=\int \frac{e^{x}(e^xdx)}{1+e^x}=\int\frac{(u-1)du}{u}=...\]

OpenStudy (anonymous):

i get it , thank you :)

OpenStudy (callisto):

Welcome :)